# Explaining Time Series Classifiers with PHAR: Rule Extraction and Fusion from Post-hoc Attributions

### v3

@@ 3.3.3 Resolution of rules ambiguity
Several equally valid rules can be derived for a single instance and require subsequent resolution. When transforming numeric explanations into symbolic rules, different valid rules may arise depending on training data and hyperparameters. Consider three salient timestep–channel pairs $t_{j_{1}}c_{j_{1}}$, $t_{j_{2}}c_{j_{2}}$, and $t_{j_{3}}c_{j_{3}}$. The following interval rules for an arbitrary instance $i$ might be derived:
@@
Here, each $t_{j_{m}}c_{j_{m}}$ denotes "channel" $c_{j_{m}}$ (in multivariate series) at timestep $t_{j_{m}}$, each interval $(\ell_{m},u_{m}]$ has lower bound $\ell_{m}$ and upper bound $u_{m}$, and $\gamma_{m},\kappa_{m}\in[0,1]$. Although $R_{i,1}$, $R_{i,2}$, and $R_{i,3}$ yield equivalent predictions due to temporal correlations in TS, rule ambiguity arises because explainers (e.g., SHAP, LIME) combined with hyperparameter-driven optimization select different variables. This optimization, which maximizes an objective defined in Section 3.3.2, resolves ambiguity within a single explainer by choosing, for example, $R_{i,\mathrm{shap}}$ for SHAP-based attribution. However, the same procedure on LIME may produce $R_{i,\mathrm{lime}}$. To handle these divergent explanations, in Section 3.4 we propose fusing the rule set $\mathcal{R}_{i}=\{\,R_{i,\mathrm{shap}},\;R_{i,\mathrm{lime}},\;R_{i,\mathrm{anchor}}\,\}$ and selecting a single, optimized rule per instance.
@@ 3.4 Rule fusion across explainers
In rule-based systems theory [16], combining rules is called rule aggregation. Our framework calls this rule fusion, as it consolidates multiple rules into exactly one whenever possible. Unlike standard aggregation, our method also applies to a single explainer, assigning each instance a rule refined via metric optimization (e.g., Lasso method discussed later).
@@
Weighted: We define a weight $w_{m}$ for each method $m$, derived from a chosen metric (e.g., confidence). We compute a weighted presence of each feature and select only those exceeding a threshold $\tau$:
@@
Conditions with $\beta_{j}\approx 0$ are discarded, yielding sparser rules. Lasso local, Lasso for short, fits per instance; Global lasso finds conditions across all instances at once to find globally important conditions. Multiple conditions per feature are combined by union. If no rules exist for an observation in Lasso global, imputation is used. Rules are first collected per observation from different explainers (e.g., SHAP, LIME, Anchor). If no rule is assigned to an observation, the most frequent rule in the predicted class is used. If none exist, the most frequent rule overall is assigned as fallback. In Lasso local (later referred as Lasso) and other methods, observations without rules are excluded.
Best: We select the single rule that maximizes a chosen metric (e.g., confidence):
@@ 4.1 Model architecture
Figure 3: Architectural overview of the ConvLSTM1D-based model used in the experiments.
@@ 4.2.1 Computational complexity of base explainers
SHAP explanations, based on kernel SHAP, have a computational complexity of $O(M^{2})$, where $M$ is the number of input features [17]. Exact SHAP computations are exponential ($O(2^{M})$), but Kernel SHAP approximates Shapley values via weighted linear regression, which still scales quadratically with $M$  [^3] https://shap.readthedocs.io/en/latest/generated/shap.KernelExplainer.html.
LIME explanations involve sampling perturbed instances and fitting a local surrogate model (typically linear regression), resulting in complexity $O(S\cdot M)$, where $S$ is the number of samples [28]. Further details are provided in the LIME documentation  [^4] https://github.com/marcotcr/lime .
Anchor explanations involve a beam search procedure over binary feature subsets, with complexity proportional to the size of the search space and sampling budget, typically $O(B\cdot S)$ where $B$ is beam width and $S$ sampling per candidate [29], see appropriate github repository [^5] https://github.com/marcotcr/anchor.
@@ 4.2.2 Computational complexity of numeric-to-rule transformation
Hyperparameter optimization guides the selection of optimal threshold and perturbation settings. The search tunes four key parameters: the percentile threshold $p\in[50,99]$, a binary flag $\emph{global}\in\{\text{true},\text{false}\}$, perturbation scale $\sigma\in[0.01,1.0]$, and number of perturbation samples $S_{p}\in[1000,10000]$ (see Table 1, Appendix). We use Optuna, a Bayesian optimization framework that adaptively explores this space and prunes unpromising trials early [2]. In the worst-case scenario: if no pruning occurs—the cost scales with the number of evaluated trials $T$, which can be as high as the total number of possible combinations in the Cartesian product of parameter ranges. However, due to pruning and adaptive sampling, the actual number of full evaluations is observed to be a small fraction of $T$. For practical use, PHAR allows bypassing this stage entirely by applying default settings, e.g., $p=90$, $\emph{global}=\text{true}$, making the optimization step optional.
@@
Interval derivation constitutes the most computationally intensive step, with complexity scaling as $O(N\cdot F^{\prime}\cdot S_{p})$, where $F^{\prime}\ll F$ denotes the average number of important features selected after thresholding, and $S_{p}$ is the number of perturbation samples generated per instance. In practice, $S_{p}$ is user-defined (typically $\leq 10000$), and both the perturbation sampling and subsequent model evaluations are highly parallelizable. This keeps the overall runtime feasible for time series datasets of moderate size, such as those commonly found in the UCR/UEA archives, where $N$ usually ranges from several hundred to a few thousand instances, and $F$ corresponds to tens or low hundreds of input features [9, 6]. This design ensures that PHAR remains computationally tractable on standard multi-core hardware.
@@ 4.2.3 Computational complexity of rule fusion
Intersection, Union, and Weighted methods select the final rule for each instance by aggregating multiple candidate rules originating from different explainability sources. These methods rely on precomputed metrics embedded within the candidate rules (e.g., confidence, coverage) and do not require re-evaluation of the predictive model. Their computational cost scales linearly with the number of rule sources $M$ and the number of explained instances $N$, resulting in an overall complexity of $O(N\cdot M)$. Since the number of candidate rule sources per instance is typically low (e.g., $M=2$ or $3$), this step introduces negligible overhead relative to earlier phases.
Lasso involves solving a sparse linear regression problem, scaling as $O(N\cdot F_{r}^{2})$ for each instance or globally across the dataset. However, the rule matrix is highly sparse, allowing efficient optimization via coordinate descent [36]. Overall, fusion contributes marginally to total computation time, with the primary bottleneck being the initial attribution computation.
@@ 4.3 Evaluation metrics and statistical testing
We compared the baseline explainers and fusions of two or three explainers. $\bar{M}$ was applied as described in Section 3.3.2. The metric $ER$ counts instances where the explainer generated rules in proportion to all cases. $\overline{\text{CONF}}$ and $\overline{\text{COV}}$, as in Section 3.3.2, were averaged over all instances. $\overline{F(n)}$ and $Med(F(n))$ represent the mean and median number of features in generated rules (for instances with rules). Furthermore, we defined composite metrics $\overline{\text{CONF}}\cdot\text{ER}$ and $\overline{\text{CONF}}\cdot\overline{\text{COV}}\cdot ER$, which combine confidence, coverage, and rule generation rates. To compare explainers and fusion methods statistical evaluation uses non-parametric Friedman test and post-hoc analysis with Nemenyi tests. Metrics are visualized for univariate and multivariate datasets to highlight key differences. Package scikit-posthocs [35] was used.
@@ 4.4 Setup: time series, explainers, and rule fusion
The datasets used in this study come from the UCR/UEA Time Series Classification Archive. Some multivariate datasets (e.g., Libras, PenDigits) originate from the UCI ML Repository but are used via their curated UCR/UEA [^6] UCR/UEA archive: https://www.cs.ucr.edu/~eamonn/time_series_data_2018/ and https://www.timeseriesclassification.com/. Selected datasets (e.g., PenDigits) originate from UCI: https://archive.ics.uci.edu/. versions [9, 6]. Our code is available at GitHub [^7] https://github.com/mozo64/papers/tree/main/phar/notebooks in the following notebooks: UCR-explainers-lime-shap-anchor.ipynb, UCR-num2rules-optuna.ipynb, UCR-fusions-results.ipynb.
We evaluated our methods on 43 datasets, including 81.4% univariate and 18.6% multivariate TS. These datasets cover diverse domains. For each dataset, a TensorFlow model classified TS samples, followed by explanations generated using Anchor, SHAP, and LIME at the variable (>1>1 for multivariate) and time point levels. SHAP and LIME were converted into rule-based explanations using the process in 3.3.2. A separate test set, unseen during training, was used for evaluation. Table A.1.2 summarizes datasets, including the number of features used in each dataset, explainers, fusion methods and resulting fused rule sets.
Anchor, LIME, and SHAP were primary baselines, forming the foundation for fused rule sets. Anchor was used in 58.1%, LIME in 65.1%, and SHAP in 62.8% of datasets. Fusion techniques (Lasso, Weighted) refined these explanations. Lasso, Lasso global, and weighted were applied to all 43 datasets (100.0%), while Best, Intersection, and Union were used in 62.8%. The latter methods were omitted when applied to a single baseline explainer since they would replicate its results. In contrast, Lasso and Weighted optimized rules even for individual explainers. As expected, the Baseline method appeared in all datasets (100.0%) as a reference. Fused rule sets combining two or three explainers were applied in 62.8% and 23.3% of datasets, respectively.
@@ 4.5.1 Pairwise and aggregated comparisons
Performance of baseline explainers: We evaluated Anchor and rules derived from LIME and SHAP under the baseline condition, representing their original, untransformed explanations. Unlike fusion methods that optimize rule structures, baselines reflect each explainer’s native performance. To compare them, we applied the Wilcoxon signed-rank test on matched datasets, ensuring robustness for non-normal data. Table 3 presents the statistical results, while Figure 8 visualizes the distribution of key metrics.
@@
Visualization of comparisons: Figure 9 visualizes metric distributions, contrasting single explainers (orange) with fusion methods (blue) for $\bar{M}$, $\overline{\text{CONF}}\cdot ER$, $\overline{\text{COV}}$, and $F(n)$. Each boxplot includes mean, standard deviation, and median values. Additional plots detail results contrasting single explainers baseline with fusion explainers: Figure 10 ($\bar{M}$), Figure 11 ($\overline{\text{CONF}}\cdot ER$), Figure 12 ($\overline{\text{COV}}$), and Figure 13 ($F(n)$).
@@ 4.5.2 Ranking-based and post-hoc analysis
Critical Difference Diagrams: CD diagrams rank methods across datasets based on their average performance. A lower rank generally indicates better results, as it corresponds to higher values of the evaluated metric, except for $\bar{F(n)}$, where lower values are preferred, making higher ranks more desirable. Horizontal bars connect methods that do not differ significantly at $\alpha=0.05$, while gaps denote statistically significant differences. Figure 4 presents CD diagrams for different fusion methods and baselines, ranking them according to their effectiveness across multiple metrics. Figure 5 extends the comparison to fused rule sets, showing how different explainers contribute to overall performance when used in combination. Additional figures in the appendix (Figures 14a14c) provide detailed comparisons for $\bar{M}$ , $ER$ , and $\bar{F(n)}$ , incorporating both explainer-level analysis (including individual explainers and their combinations) and specific fusion methods applied to these rule sets.
@@ 4.6 Summary of evaluation results
Anchor consistently performed best in $\bar{M}$ and minimizing feature count, but was not dominant in $\overline{\text{CONF}}$, $\overline{\text{COV}}$, and $ER$. Among explainers, SHAP outperformed LIME in $\overline{\text{CONF}}$, and Anchor+SHAP surpassed Anchor. Most multi-explainer fusions improved $\overline{\text{COV}}$, while no significant differences were found for $ER$. Fusion methods had mixed effectiveness and choosing the right method is crucial. In compound metrics, Lasso, Weighted, and Best consistently improved performance. Some methods, such as Union, increased feature count significantly, while Lasso and Intersection minimized it. CD Diagrams confirmed that fusions often outperformed individual explainers. The best methods depended on the metric: Lasso, Weighted, and Best excelled in $\bar{M}$ and compound metrics, while Lasso global was particularly effective in $\overline{\text{COV}}$. Among explainer combinations, Anchor-based fusions ranked highest, but LIME+SHAP also performed well in $\overline{\text{CONF}}$ and $ER$.

### v4

@@ 3.3.3 Resolution of rules ambiguity
Several equally valid rules can be derived for a single instance and require subsequent resolution. When transforming numeric explanations into symbolic rules, different valid rules may arise depending on training data and hyperparameters. Consider three distinct timestep–channel pairs $t_{j_{1}}c_{j_{1}}$, $t_{j_{2}}c_{j_{2}}$, and $t_{j_{3}}c_{j_{3}}$. The following interval rules might be derived for an arbitrary instance $i$:
@@
Here, each $t_{j_{m}}c_{j_{m}}$ denotes "channel" $c_{j_{m}}$ (in multivariate series) at timestep $t_{j_{m}}$, each interval $(\ell_{m},u_{m}]$ has a lower bound $\ell_{m}$ and an upper bound $u_{m}$, and $\gamma_{m},\kappa_{m}\in[0,1]$. Although $R_{i,1}$, $R_{i,2}$, and $R_{i,3}$ yield equivalent predictions due to temporal correlations in TS, rule ambiguity arises because explainers (e.g. SHAP, LIME) combined with hyperparameter-driven optimization select different variables. This optimization, which maximizes an objective defined in Section 3.3.2, resolves the ambiguity within a single explainer by choosing, for example, $R_{i,\mathrm{shap}}$ for SHAP-based-based attribution. However, the same procedure on LIME may produce $R_{i,\mathrm{lime}}$. To handle these divergent explanations, in Section 3.4 we propose fusing the rule set $\mathcal{R}_{i}=\{\,R_{i,\mathrm{shap}},\;R_{i,\mathrm{lime}},\;R_{i,\mathrm{anchor}}\,\}$ and selecting a single optimized rule per instance.
@@ 3.4 Rule fusion across explainers
In rule-based systems theory [ligeza_rule_based], combining rules is called rule aggregation. Our framework calls this rule fusion, as it consolidates multiple rules into exactly one whenever possible. Unlike standard aggregation, our method also applies to a single explainer, assigning each instance a rule refined via metric optimization (e.g. Lasso method discussed later).
@@
Weighted: We define a weight $w_{m}$ for each method $m$, derived from a chosen metric (e.g. confidence). We compute a weighted presence of each feature and select only those exceeding a threshold $\tau$:
@@
Conditions with $\beta_{j}\approx 0$ are discarded, yielding sparser rules. Lasso local, Lasso for short, fits per instance; Global lasso finds conditions across all instances at once to find globally important conditions. Multiple conditions per feature are combined by union. If no rules exist for an observation in Lasso global, imputation is used. Rules are first collected per observation from different explainers (e.g. SHAP, LIME, Anchor). If no rule is assigned to an observation, the most frequent rule in the predicted class is used. If none exist, the most frequent rule overall is assigned as fallback. In Lasso local (later referred as Lasso) and other methods, observations without rules are excluded.
Best: We select the single rule that maximizes a chosen metric (e.g. confidence):
@@ 4.1 Model architecture
Figure 3: Overview of the architecture of the ConvLSTM1D-based model used in the experiments.
@@ 4.2.1 Computational complexity of base explainers
SHAP explanations, based on kernel SHAP, have a computational complexity of $O(M^{2})$, where $M$ is the number of input features [lundberg2017unifiedapproachinterpretingmodel]. Exact SHAP computations are exponential ($O(2^{M})$), but Kernel SHAP approximates Shapley values via weighted linear regression, which still scales quadratically with $M$  [^3] https://shap.readthedocs.io/en/latest/generated/shap.KernelExplainer.html.
LIME explanations involve sampling perturbed instances and fitting a local surrogate model (typically linear regression), resulting in complexity $O(S\cdot M)$, where $S$ is the number of samples [ribeiro2016why]. Further details are provided in the LIME documentation  [^4] https://github.com/marcotcr/lime .
Anchor explanations involve a beam search procedure over binary feature subsets, with complexity proportional to the size of the search space and sampling budget, typically $O(B\cdot S)$ where $B$ is beam width and $S$ sampling per candidate [Ribeiro_Singh_Guestrin_2018], see appropriate github repository [^5] https://github.com/marcotcr/anchor.
@@ 4.2.2 Computational complexity of numeric-to-rule transformation
Hyperparameter optimization guides the selection of optimal threshold and perturbation settings. The search tunes four key parameters: the percentile threshold $p\in[50,99]$, a binary flag $\emph{global}\in\{\text{true},\text{false}\}$, perturbation scale $\sigma\in[0.01,1.0]$, and number of perturbation samples $S_{p}\in[1000,10000]$ (see Table 1, Appendix). We use Optuna, a Bayesian optimization framework that adaptively explores this space and prunes unpromising trials early [optuna_2019]. In the worst-case scenario: if no pruning occurs—the cost scales with the number of evaluated trials $T$, which can be as high as the total number of possible combinations in the Cartesian product of parameter ranges. However, due to pruning and adaptive sampling, the actual number of full evaluations is observed to be a small fraction of $T$. For practical use, PHAR allows bypassing this stage entirely by applying default settings, e.g. $p=90$, $\emph{global}=\text{true}$, making the optimization step optional.
@@
Interval derivation constitutes the most computationally intensive step, with complexity scaling as $O(N\cdot F^{\prime}\cdot S_{p})$, where $F^{\prime}\ll F$ denotes the average number of important features selected after thresholding, and $S_{p}$ is the number of perturbation samples generated per instance. In practice, $S_{p}$ is user-defined (typically $\leq 10000$), and both the perturbation sampling and subsequent model evaluations are highly parallelizable. This keeps the overall runtime feasible for time series datasets of moderate size, such as those commonly found in the UCR/UEA archives, where $N$ usually ranges from several hundred to a few thousand instances, and $F$ corresponds to tens or low hundreds of input features [Dau2019UCRArchive, Bagnall2018UEAMultivariate]. This design ensures that PHAR remains computationally tractable on standard multi-core hardware.
@@ 4.2.3 Computational complexity of rule fusion
Intersection, Union, and Weighted methods select the final rule for each instance by aggregating multiple candidate rules originating from different explainability sources. These methods rely on precomputed metrics embedded within the candidate rules (e.g. confidence, coverage) and do not require re-evaluation of the predictive model. Their computational cost scales linearly with the number of rule sources $M$ and the number of explained instances $N$, resulting in an overall complexity of $O(N\cdot M)$. Since the number of candidate rule sources per instance is typically low (e.g. $M=2$ or $3$), this step introduces negligible overhead relative to earlier phases.
Lasso involves solving a sparse linear regression problem, scaling as $O(N\cdot F_{r}^{2})$ for each instance or globally across the dataset. However, the rule matrix is highly sparse, allowing efficient optimization via coordinate descent [Tibshirani1996].
Overall, fusion contributes marginally to total computation time, with the primary bottleneck being the initial attribution computation.
@@ 4.3 Evaluation metrics and statistical testing
We compared the baseline explainers and fusions of two or three explainers. $\bar{M}$ was applied as described in Section 3.3.2. The metric $ER$ counts instances where the explainer generated rules in proportion to all cases. $\overline{\text{CONF}}$ and $\overline{\text{COV}}$, as in Section 3.3.2, were averaged over all instances. $\overline{F(n)}$ and $Med(F(n))$ represent the mean and median number of features in generated rules (for instances with rules). Furthermore, we defined composite metrics $\overline{\text{CONF}}\cdot\text{ER}$ and $\overline{\text{CONF}}\cdot\overline{\text{COV}}\cdot ER$, which combine confidence, coverage, and rule generation rates. To compare explainers and fusion methods statistical evaluation uses non-parametric Friedman test and post-hoc analysis with Nemenyi tests. Metrics are visualized for univariate and multivariate datasets to highlight key differences. Package scikit-posthocs [Terpilowski2019] was used.
@@ 4.4 Setup: time series, explainers, and rule fusion
The datasets used in this study come from the UCR/UEA Time Series Classification Archive. Some multivariate datasets (e.g. Libras, PenDigits) originate from the UCI ML Repository but are used via their curated UCR/UEA [^6] UCR/UEA archive: https://www.cs.ucr.edu/~eamonn/time_series_data_2018/ and https://www.timeseriesclassification.com/. Selected datasets (e.g. PenDigits) originate from UCI: https://archive.ics.uci.edu/. versions [Dau2019UCRArchive, Bagnall2018UEAMultivariate]. Our code is available at GitHub [^7] https://github.com/mozo64/papers/tree/main/phar/notebooks in the following notebooks: UCR-explainers-lime-shap-anchor.ipynb, UCR-num2rules-optuna.ipynb, UCR-fusions-results.ipynb.
We evaluated our methods on 43 datasets, including 81.4% univariate and 18.6% multivariate TS. These datasets cover diverse domains. For each dataset, a TensorFlow model classified the TS samples, followed by explanations generated using Anchor, SHAP, and LIME at the variable (>1>1 for multivariate) and time point levels. SHAP and LIME were converted into rule-based explanations using the process in 3.3.2. A separate test set, unseen during training, was used for evaluation. Table A.1.2 summarizes the datasets, including the number of features used in each dataset, explainers, fusion methods, and resulting fused rule sets.
Across the benchmark, Anchor, LIME, and SHAP served as primary baselines and provided the rule sets on which fusion methods were built. Due to the computational cost of running some explainers and fusion procedures on long, high-dimensional time series, not every explainer could be executed on all 43 datasets within our runtime budget. Consequently, Anchor, LIME, and SHAP contributed baseline rules in 58.1%, 65.1%, and 62.8% of the datasets, respectively.
Fusion techniques such as Lasso, Lasso global, and Weighted were applied to all datasets (100.0%), since they can refine rules even when only a single explainer is available. In contrast, Best, Intersection, and Union were only used when at least two distinct baseline explainers were available, otherwise, they would trivially replicate the underlying baseline. As a result, these three fusion methods were applied in 62.8% of the datasets. By construction, the Baseline method appears in all datasets (100.0%) as a reference. Finally, fused rule sets combining two and three explainers were applicable in 62.8% and 23.3% of the datasets, respectively, reflecting the fraction of problems for which two or all three baseline explainers were computed.
@@ 4.5.1 Pairwise and aggregated comparisons
Performance of baseline explainers: We evaluated Anchor and rules derived from LIME and SHAP under the baseline condition, representing their original, untransformed explanations. Unlike fusion methods that optimize rule structures, baselines reflect each explainer’s native performance. To compare them, we applied the Wilcoxon signed-rank test on matched datasets, ensuring robustness for non-normal data. Table 3 presents the statistical results, while Figure A1 visualizes the distribution of key metrics.
@@
Visualization of comparisons: Figure A2 visualizes metric distributions, contrasting single explainers (orange) with fusion methods (blue) for $\bar{M}$, $\overline{\text{CONF}}\cdot ER$, $\overline{\text{COV}}$, and $F(n)$. Each boxplot includes mean, standard deviation, and median values. Additional plots detail results contrasting single explainers baseline with fusion explainers: Figure A3 ($\bar{M}$), Figure A4 ($\overline{\text{CONF}}\cdot ER$), Figure A5 ($\overline{\text{COV}}$), and Figure A6 ($F(n)$).
@@ 4.5.2 Ranking-based and post-hoc analysis
Critical Difference Diagrams (CDD): CDD rank methods across datasets based on their average performance. A lower rank generally indicates better results, as it corresponds to higher values of the evaluated metric, except for $\bar{F(n)}$, where lower values are preferred, making higher ranks more desirable. Horizontal bars connect methods that do not differ significantly at $\alpha=0.05$, while gaps denote statistically significant differences. Figure 4 presents CDD for different fusion methods and baselines, ranking them according to their effectiveness across multiple metrics. Figure 5 extends the comparison to fused rule sets, showing how different explainers contribute to overall performance when used in combination. Additional figures in the appendix (Figures 14a14c) provide detailed comparisons for $\bar{M}$ , $ER$ , and $\bar{F(n)}$ , incorporating both explainer-level analysis (including individual explainers and their combinations) and specific fusion methods applied to these rule sets.
@@ 4.6 Summary of evaluation results
Anchor consistently performed best in $\bar{M}$ and minimizing feature count, but was not dominant in $\overline{\text{CONF}}$, $\overline{\text{COV}}$, and $ER$. Among explainers, SHAP outperformed LIME in $\overline{\text{CONF}}$, and Anchor+SHAP surpassed Anchor. Most multi-explainer fusions improved $\overline{\text{COV}}$, while no significant differences were found for $ER$. Fusion methods had mixed effectiveness and choosing the right method is crucial. In compound metrics, Lasso, Weighted, and Best consistently improved performance. Some methods, such as Union, increased feature count significantly, while Lasso and Intersection minimized it. CDD confirmed that fusions often outperformed individual explainers. The best methods depended on the metric: Lasso, Weighted, and Best excelled in $\bar{M}$ and compound metrics, while Lasso global was particularly effective in $\overline{\text{COV}}$. Among explainer combinations, Anchor-based fusions ranked highest, but LIME+SHAP also performed well in $\overline{\text{CONF}}$ and $ER$.